y varies directly as x and inversely as the square of z. y=38 when x=32 and z=4. Find y when x=75 and z=5.
y=

Answer by Cromlix(4381)

y varies directly as x and inversely as the square of z
y = kx/z^2
(k is the constant of variation)
To find k use all the values given:-
y = kx/z^2
38 = k * 32/4^2
38 = k32/16
38 = 2k
k = 19
Formula => y = 19x/z^2
Solve:
y = 19x/z^2
y = 19*75/5^2
y = 1425/25
y = 57
Hope this helps
:-)